OpenStudy (anonymous):

( 4 * 9 ^ x ) - ( 7 * 12 ^ x ) - ( 3 * 16 ^ x ) = 0

OpenStudy (anonymous):

Sorry its ( 4 * 9 ^ x ) - ( 7 * 12 ^ x ) + ( 3 * 16 ^ x ) = 0

OpenStudy (anonymous):

You can check that x =0 is a solution.

OpenStudy (lucaz):

and x=1

OpenStudy (anonymous):

Also x=1

OpenStudy (anonymous):

I was searching for something like more precise. For example we have 4 * 3 ^ 2x - 7 * 4 ^ x * 3 ^ x + 3 * 4 ^ 2x = 0 and if we replace 3 ^ x with i and 4 ^ x with j we come to this: 4 * ( i ^ 2 ) - 7 * i * j + 3 * j ^ 2 = 0 and then...

OpenStudy (anonymous):

\[ 4\ 9^x-7\ 12^x+3\ 16^x=\left(4\ 3^x-3\ 4^x\right) \left(3^x-4^x\right)=0 \]

OpenStudy (anonymous):

x=0 and x=1 are the only solutions. Look at http://www.wolframalpha.com/input/?i=plot+4+9^x+-+7+12^x+%2B+3+16^x+from+x+%3D-2+to+1.2